# Higher order linear differential equations for unitary matrix integrals: applications and generalisations

### v1

@@ 1.1 A unitary matrix integral and the longest increasing subsequence problem
the same average with a further factor of $(\det U)^{q}$, and their $\beta$ generalisation. In (1.1) $U(l)$ refers to the group of $l\times l$ unitary matrices sampled with Haar measure. This matrix integral first appeared in the theoretical physics literature on two-dimensional Yang-Mills gauge theory as a partition function [39]. However it is not in this context that interest in differential equations for matrix integrals first arose. Rather, it came about due to the realisation that the coefficients $\{T_{l}(N)\}$ say of $z^{2N}/(N!)^{2}$ in the power series of (1.1) (note that this is even in $z$) are for a given $l$ an integer family, e.g. for $l=4$ and starting at $N=1$ [40, sequence A047889]
@@
Thus, according to [42], $T_{l}(N)$ counts the number or permutations of $\{1,2,\dots,N\}$ such that the length of the longest increasing subsequence is less than or equal to $l$ (this number is $N!$ for $N\leq l$ as illustrated in (1.2)).
@@
where $I_{\nu}(x)={1\over 2\pi}\int_{0}^{2\pi}e^{x\cos\theta}e^{i\nu\theta}\,d\theta$, $\nu\in\mathbb{Z}$, is the modified Bessel function of the first kind (i.e. I-Bessel function), although $T_{l}(N)$ was defined as the number of pairs of standard tableaus of content $N$ with length of the first row less than or equal to $l$. For an explanation of these notions, and the equivalence to increasing subsequences of permutations, see e.g. [31, 43]. The reason why the Toeplitz determinant in (1.4) is equivalent to the matrix integral in (1.3) follows from the general identities
@@
Nonconstructive arguments were given in [32] that the generating function in (1.3) is $D$-finite, meaning that it must satisfy a linear differential equation with polynomial coefficients. Later, in [9], these differential equations where determined explicitly for $l\leq 7$. For example, in the case $l=4$ and thus relating to (1.2) one has that (1.3) with $z\mapsto\sqrt{z}$ is a solution of the fifth order equation [9, last collection of displayed equations]
@@
(there is also a known linear differential equation of degree four for the power series formed according to the rule that the coefficient of $z^{k}$ is given by the $k$-th member of (1.2), and the coefficient of $z^{0}$ is unity [13, §3.3]). In the present paper, using the theory of Selberg correlation integrals [16], [2], [20, §4.6], [28], we will show how to systematically obtain a first order $(l+1)\times(l+1)$ matrix linear differential equation, which has as the first component of its vector solution equal to (1.3). As a preliminary step, the matrix average is expressed in terms of a certain Jack polynomial based hypergeometric function of $l$ variables, specialised to the case that all the variables are equal; see Section 2. For a closely related class of hypergeometric functions, earlier work [25] has identified a matrix differential equation characterisation. It is shown in Section 3 how the Jack polynomial based hypergeometric functions of present interest can be deduced as confluent limits of those appearing in [25], allowing for these too to be similarly characterised. This implies a scalar linear differential equation of degree $l+1$ satisfied by (1.3) (Proposition 4.1 below), which with the use of computer algebra can be made explicit. As an illustration, we do this explicitly for $l=8$, which is the first case beyond the listing in [9]; see Proposition 4.2 below.
@@
Motivating taking up the problem of linear differential equation equations satisfied by (1.1) at the present time, and furthermore motivating a more extensive study, is the primary role played by the modification of (1.1) and the Toeplitz determinant in (1.4)
@@
in the recent studies [35], [36]. Here the second equality views the determinant in the Hankel class (i.e. indices a function of $j+k$) which is the structure emphasised in these studies. The theme of [35], [36] is the question of statistical properties, particularly moments, of higher order derivatives of the Riemann zeta function on the critical line. Fundamentally, in keeping with the celebrated work [34], it is hypothesised that these zeta function moments coincide with the scaled limiting moments of the corresponding derivative of the characteristic polynomial of a Haar distributed random unitary matrix; see [33, 15, 17, 47, 8, 7, 3, 6, 23] for earlier work in this direction, and [4, 44, 1, 5] for subsequent developments. In particular, in relation to the second derivative, this leads to the study of [35]
@@
In [36], certain reasoning is given that leads to the conclusion that like (1.1) $\tau_{l}(x)$ satisfies a linear differential equation of degree $l+1$ (see Appendix A for an overview). As mentioned in the opening paragraph, our approach to systematically deriving a linear differential equation that (1.1) satisfies allows for an extra factor $(\det U)^{q}$ to be included for general $q$, and thus includes (1.7). Thus in this regards our work complements [36]. We consider this further in §4.3, and discuss too the advantage of the linear differential equation characterisation over an alternative nonlinear differential equation characterisation put forward in [30].
@@ 2 Jack polynomial based hypergeometric function form
It has been revised in the Introduction that the matrix integral (1.1) has the interpretations as a partition function for a two-dimensional gauge theory, and as the generating function for counting longest increasing subsequences. There is also an occurrence of (1.1) as a limiting gap probability at the hard edge of a positive definite matrix ensemble with unitary symmetry. To specify this, define complex Wishart matrices $W=X^{\dagger}X$ with $X$ an $n\times N$ $(n\geq N)$ standard complex Gaussian matrix (also referred to as a rectangular GinUE matrix [14]). Let $E_{N}((0,s);a)$ with l=nNl=n-N denote the probability of no eigenvalues in $(0,s)$. In [19] it was established that
$E_{\beta=2}^{\rm hard}((0,s);l):=\lim_{N\to\infty}E_{N}((0,s/4N);l)=e^{-s/4}\Big{\langle}e^{(s^{1/2}/2){\rm Tr}(U+U^{\dagger})}\Big{\rangle}_{U(l)}.$ (2.1)
The most significant contribution in [19] for present purposes was the identification of the RHS of (2.1) (and generalisations) as a Jack polynomial based generalised hypergeometric function.
To specify this class of multivariable special function, we first recall that, with $\mathbf{x}=(x_{1},\dots,x_{n})$ the Jack polynomials $P_{\kappa}^{(\alpha)}(\mathbf{x})$ are symmetric polynomials of $n$ variables [38, Ch. VI.6], [20, Ch. 12], [37, Ch. 7]. They are dependent on a parameter $\alpha>0$ and labelled by a partition $\kappa:=(\kappa_{1},\dots,\kappa_{n})$; in the latter each part $\kappa_{i}\in\mathbb{Z}_{\geq 0}$ is ordered as $\kappa_{1}\geq\kappa_{2}\geq\cdots\geq\kappa_{n}$. These polynomials are eigenfunctions of the differential operator
@@
$P_{\kappa}^{(\alpha)}(\mathbf{x})\mapsto\alpha^{|\kappa|}P_{\kappa}^{(\alpha)}(\mathbf{x})/h_{\kappa}^{\prime},$
where the precise definition of $h_{\kappa}^{\prime}$ (which depends too on $\alpha$) can be found in e.g. [20, Eq. (12.46)]. Playing the role of the Pochhammer symbol is the quantity
@@
see e.g. [20, Eq. (12.46)]. With this notation we then have as the series definition of the generalised hypergeometric function in question (see e.g. [20, §13.1])
@@
(see too [20, Eq. (13.27)] and [24, Eq. (4.32)]), where $c\in\mathbb{Z}_{\geq 1}$ (this restriction can be removed by suitably deforming the implied circular contours in (2) to Hankel loops [21, Proof of Prop. 2]) and
@@
for a derivation of this result, and its generalisation for $\alpha=1/2$ and $\alpha=2$, based on the theory of zonal polynomials, see [27]). Extending this, for any $q\in\mathbb{Z}_{\geq 0}$ we read off that
@@ 3 Matrix differential equation characterisation
The significance of this is that for $\mathbf{x}=(t)^{n}$, since the 2022 work [25] we have available a matrix differential equation characterisation of ${\vphantom{F}}_{1}^{\mathstrut}F_{1}^{(\alpha)}(a;b;\mathbf{x})$. This comes about in the setting of the integral representation [20, Cor. 2.1]
@@
valid for Re(a)>1(a)>-1, Re(b)>1(b)>-1. Here $S_{n}(a,b,\beta)$ is the Selberg integral (see e.g. [20, Ch. 4]), which normalises the RHS so that it is equal to unity for $s=0$. The matrix differential equation, which we will first present in the form of a differential-difference recurrence, relates to the family of multi-dimensional integrals
$\hat{H}_{p}(x)={1\over C_{p}^{n}}\int_{0}^{1}dt_{1}\cdots\int_{0}^{1}dt_{n}\,\prod_{l=1}^{n}t_{l}^{a}(1-t_{l})^{b-1}e^{-x\sum_{l=1}^{n}t_{l}}\\ \times\prod_{1\leq j<k\leq n}|t_{k}-t_{j}|^{\beta}e_{p}(1-t_{1},\dots,1-t_{n}),$ (3.3)
@@
[25, Cor. 2.5] The multidimesional integrals $\{\hat{H}_{p}(x)\}_{p=0}^{n}$ satisfy the differential-difference system
$(n-p)x\hat{H}_{p+1}(x)=\Big{(}(n-p)x+\tilde{B}_{p}\Big{)}\hat{H}_{p}(x)+x{d\over dx}\hat{H}_{p}(x)-\tilde{D}_{p}\hat{H}_{p-1}(x),\quad(p=0,\dots,n),$ (3.4)
@@
$\tilde{B}_{p}=p\Big{(}a+b+1+(\beta/2)(2n-p-1)\Big{)},\qquad\tilde{D}_{p}=p\Big{(}(\beta/2)(n-p)+b\Big{)}.$
@@
Introduce the column vector $\hat{\mathbf{H}}(x)=[\hat{H}_{p}(x)]_{p=0}^{n}$, and the bi-diagonal matrices
@@
$\mathbf{Y}$ $=-{\rm diag}\,[\tilde{B}_{0},\tilde{B}_{1},\dots,\tilde{B}_{n}]+{\rm diag}^{-}\,[\tilde{D}_{1},\tilde{D}_{2},\dots,\tilde{D}_{n}].$ (3.5)
@@
$x{d\over dx}\hat{\mathbf{H}}(x)=(x\mathbf{X}+\mathbf{Y})\hat{\mathbf{H}}(x).$ (3.6)
As a consequence [25, Remark 4.19], introducing the vector of power series solutions
$\hat{\mathbf{H}}(x)=\Big{[}\sum_{k=0}^{\infty}c_{p,k}x^{k}\Big{]}_{p=0}^{n},$
@@
$c_{0,0}=S_{n}(a,b-1,\beta),\quad c_{n,0}=S_{n}(a,b,\beta),\quad c_{n-p,0}=c_{n,0}\prod_{s=1}^{p}{\tilde{B}_{n+1-s}\over\tilde{D}_{n+1-s}},\>\>(p=1,\dots,n)$
@@
We now want to make use of (3.1) in Proposition 3.1, in view of both the first and last components of $\hat{\mathbf{H}}(x)$ relating to $_{1}F^{(2/\beta)}_{1}$.
@@
Consider the differential-difference system
$(n-p)x\hat{I}_{p+1}(x)=\tilde{B}_{p}\hat{I}_{p}(x)+x{d\over dx}\hat{I}_{p}(x)-p\hat{I}_{p-1}(x),\quad(p=0,\dots,n),$ (3.8)
where $\tilde{B}_{p}$ is as in (3.4). Introduce the column vector $\hat{\mathbf{I}}(x)=[\hat{I}_{p}(x)]_{p=0}^{n}$, and the matrices
$\tilde{\mathbf{X}}={\rm diag}^{+}\,[n,n-1,\dots,1],\quad\tilde{\mathbf{Y}}=-{\rm diag}\,[\tilde{B}_{0},\tilde{B}_{1},\dots,\tilde{B}_{n}]+{\rm diag}^{-}\,[1,2,\dots,n].$ (3.9)
@@
$x{d\over dx}\hat{\mathbf{I}}(x)=(x\tilde{\mathbf{X}}+\tilde{\mathbf{Y}})\hat{\mathbf{I}}(x).$ (3.10)
@@
$\hat{I}_{0}(x)$ $={\vphantom{F}}_{0}^{\mathstrut}F_{1}^{(2/\beta)}\Big{(}\beta(n-1)+a+b+1;(x)^{n}\Big{)},$ (3.12)
$\hat{I}_{n}(x)$ $=d_{n,0}\,{\vphantom{F}}_{0}^{\mathstrut}F_{1}^{(2/\beta)}\Big{(}\beta(n-1)+a+b+2;(x)^{n}\Big{)},$ (3.13)
and, provided all the $\tilde{B}_{s}$ are nonzero,
$d_{p,0}={p!\over\prod_{s=1}^{p}\tilde{B}_{s}},\>\>(p=1,\dots,n).$ (3.14)
@@
$(k{\mathbf{I}}_{n+1}-\tilde{\mathbf{Y}})\mathbf{d}_{k}=\tilde{\mathbf{X}}\mathbf{d}_{k-1},\>\>(k=1,2,\dots).$ (3.15)
@@
To deduce (3.8) and (3.10), we substitute in (3.4) and (3.6), respectively, $x\mapsto-x/a$, $\hat{H}_{p}(-x/a)=b^{p}\hat{I}_{p}(x)$ and take the limit $a\to\infty$ with $a+b$ fixed. The evaluations (3.13) follow from the evaluations of $\hat{H}_{n}(x)$, $\hat{H}_{0}(x)$ in terms of $\vphantom{F}{}_{1}^{\mathstrut}F_{1}^{(2/\beta)}$ noted in the sentence above Proposition 3.1. The evaluation (3.14) can be calculated as the eigenvector of $\tilde{\mathbf{Y}}$ corresponding to the eigenvalue 0, normalised to have its first component equal to unity. The vector recurrence relation (3.15) follows by substituting (3.11) in (3.10) and equating like powers of $x$. ∎
[…]
$\tilde{B}_{p}=p\Big{(}q+1+(\beta/2)(l-p)\Big{)},$ (3.16)
@@
$\mathbf{d}_{k}=(k{\mathbf{I}}_{l+1}-\tilde{\mathbf{Y}})^{-1}\tilde{\mathbf{X}}\mathbf{d}_{k-1},\>\>(k=1,2,\dots)$ (3.17)
@@
$\mathbf{R}_{N,l}:=(N{\mathbf{I}}_{l+1}-\tilde{\mathbf{Y}})^{-1}\tilde{\mathbf{X}}((N-1){\mathbf{I}}_{l+1}-\tilde{\mathbf{Y}})^{-1}\tilde{\mathbf{X}}\cdots({\mathbf{I}}_{l+1}-\tilde{\mathbf{Y}})^{-1}\tilde{\mathbf{X}}$.
@@
1. Suppose we replace $q$ by q1q-1 in (3.16). Then setting $q=0$ we observe that the column vector $\delta_{p,l}]_{p=0}^{l}$ is an eigenvector of $\tilde{\mathbf{Y}}$ with $n=l$, allowing for the choice of initial vector corresponding to $d_{p,0}=\delta_{p,l}$ ($p=0,\dots,l$). Using this initial vector, an alternative formula to (3.18) is
@@
where the notation $(\cdot)_{l+1}$ denotes the $(l+1)$-th entry (final entry) in the corresponding column vector. In the following section, when referring to $\hat{I}_{n}(x)$ in the case $q=0$ as it relates to (2.8), it will be assumed that the initial vector has been similarly modified (and thus in (3.13) $d_{n,0}$ is replaced by unity — without this modification (3.13) is invalid since $d_{n,0}$ involves a division by zero).
2. For efficient computation using the matrix differential equation, it is advantageous to forgo (3.18) in favour of a recursive computation based on (3.15) — we take up this point further at the end of §4.2.

### v2

@@ Abstract
3 Department of Mathematics, Technical University of Munich, 80290 München, Germany.
Email: bornemann@tum.de
@@ 1.1 A unitary matrix integral and the longest increasing subsequence problem
the same average with a further factor of $(\det U)^{q}$, and their $\beta$ generalisation. In (1.1) $U(l)$ refers to the group of $l\times l$ unitary matrices sampled with Haar measure. This matrix integral first appeared in the theoretical physics literature on two-dimensional Yang-Mills gauge theory as a partition function [43]. However it is not in this context that interest in differential equations for matrix integrals first arose. Rather, it came about due to the realisation that the coefficients $\{T_{l}(N)\}$ say of $z^{2N}/(N!)^{2}$ in the power series of (1.1) (note that this is even in $z$) are for a given $l$ an integer family, e.g. for $l=4$ and starting at $N=1$ [44, sequence A047889]
@@
Thus, according to [46], $T_{l}(N)$ counts the number or permutations of $\{1,2,\dots,N\}$ such that the length of the longest increasing subsequence is less than or equal to $l$ (this number is $N!$ for $N\leq l$ as illustrated in (1.2)). An alternative specification of $T_{l}(N)$, used in e.g. [10], is as the number of pairs of standard tableaux of content $N$ with length of the first row less than or equal to $l$. For an explanation of these notions, and the equivalence of the two descriptions, see e.g. [34, 47].
@@
where $I_{\nu}(x)={1\over 2\pi}\int_{0}^{2\pi}e^{x\cos\theta}e^{i\nu\theta}\,d\theta$, $\nu\in\mathbb{Z}$, is the modified Bessel function of the first kind (i.e. I-Bessel function). The reason why the Toeplitz determinant in (1.4) is equivalent to the matrix integral in (1.3) follows from the general identities
@@
Arguments were given in [35] that the generating function in (1.3) is $D$-finite, meaning that it must satisfy a linear differential equation with polynomial coefficients; for more on this see the first subsection of Appendix B. Later, in [10], these differential equations where determined explicitly for $l\leq 7$. For example, in the case $l=4$ and thus relating to (1.2) one has that (1.3) with $z\mapsto\sqrt{z}$ is a solution of the fifth order equation [10, last collection of displayed equations]
@@
(there is also a known linear differential equation of degree four for the power series formed according to the rule that the coefficient of $z^{k}$ is given by the $k$-th member of (1.2), and the coefficient of $z^{0}$ is unity [14, §3.3]). In the present paper, using the theory of Selberg correlation integrals [19], [2], [23, §4.6], [31], we will show how to systematically obtain a first order $(l+1)\times(l+1)$ matrix linear differential equation, which has as the first component of its vector solution equal to (1.3). As a preliminary step, the matrix average is expressed in terms of a certain Jack polynomial based hypergeometric function of $l$ variables, specialised to the case that all the variables are equal; see Section 2. For a closely related class of hypergeometric functions, earlier work [28] has identified a matrix differential equation characterisation. It is shown in Section 3 how the Jack polynomial based hypergeometric functions of present interest can be deduced as confluent limits of those appearing in [28], allowing for these too to be similarly characterised. This implies a scalar linear differential equation of degree $l+1$ satisfied by (1.3) (Proposition 4.1 below), which with the use of computer algebra can be made explicit. As an illustration, we do this explicitly for $l=8$, which is the first case beyond the listing in [10]; see Proposition 4.2 below.
Efficient methods for the computation of $\{T_{l}(N)\}$ have come to the fore in the recent works [29, 12] (see too [13]), where a primary aim was to study a scaling regime in the neighbourhood of $l=2\sqrt{N}$ for $N$ large. A precise limit law for the corresponding scaled distribution has been known since [7], and these recent works set about quantifying correction terms. With the numerical tabulations achieved in these works, clear functional forms of corrections were able to exhibited graphically. Such computations were carried out using the known relation for (1.1) and a particular $\sigma$-Painlevé III equation [53, 32], and also its Chazy-I equivalent form. In the second subsection of Appendix B a comparitive discussion of the computational complexity associated with the latter characterisation, and that based on the present matrix differential equation characterisation via the vector recurrence (3.15), is given.
@@
We have a second main motivation for considering the characterisation of the matrix integral (1.1), and that of modifications of this integral, in terms of a linear differential equation equations satisfied by (1.1). This is the primary role played by the modification of (1.1) and the Toeplitz determinant in (1.4)
@@
in the recent studies [38], [39]. Here the second equality views the determinant in the Hankel class (i.e. indices a function of $j+k$), which is the structure emphasised in these studies. The theme of [38], [39] is the question of statistical properties, particularly moments, of higher order derivatives of the Riemann zeta function on the critical line. Fundamentally, in keeping with the celebrated work [37], it is hypothesised that these zeta function moments coincide with the scaled limiting moments of the corresponding derivative of the characteristic polynomial of a Haar distributed random unitary matrix; see [36, 18, 20, 55, 9, 8, 3, 6, 26] for earlier work in this direction, and [4, 50, 1, 5, 16] for subsequent developments. In particular, in relation to the second derivative, this leads to the study of [38]
@@
In [39], certain reasoning is given that leads to the conclusion that like (1.1), $\tau_{l}(x)$ satisfies a linear differential equation of degree $l+1$ (see Appendix A for an overview). As mentioned in the opening paragraph, our approach to systematically deriving a linear differential equation that (1.1) satisfies allows for an extra factor $(\det U)^{q}$ to be included for general $q$, and thus includes (1.7). Thus in this regards our work complements [39]. We consider this further in §4.3, and discuss too the advantage of the linear differential equation characterisation over an alternative nonlinear differential equation characterisation put forward in [33].
@@ 2 Jack polynomial based hypergeometric function form
It has been revised in the Introduction that the matrix integral (1.1) has the interpretations as a partition function for a two-dimensional gauge theory, and as the generating function for counting longest increasing subsequences. There is also an occurrence of (1.1) as a limiting gap probability at the hard edge (neighbourhood of the origin) of a positive definite matrix ensemble with unitary symmetry. To specify this, define complex Wishart matrices $W=X^{\dagger}X$ with $X$ an $n\times N$ $(n\geq N)$ standard complex Gaussian matrix (also referred to as a rectangular GinUE matrix [17]). Let $E_{N}((0,s);l)$ with l=nNl=n-N denote the probability of no eigenvalues in $(0,s)$ in this class of complex Wishart ensembles. In [22] it was established that
$E_{\beta=2}^{\rm hard}((0,s);l):=\lim_{N\to\infty}E_{N}((0,s/(4N));l)=e^{-s/4}\Big\langle e^{(s^{1/2}/2){\rm Tr}(U+U^{\dagger})}\Big\rangle_{U(l)}.$ (2.1)
(Here the use of the symbol $\beta$ as a subscript on the LHS is in the context of the underlying Dyson index; see (2.10) below and subsequent text.) The most significant contribution in [22] for present purposes was the identification of the RHS of (2.1) (and generalisations) as a Jack polynomial based generalised hypergeometric function.
To specify this class of multivariable special function, we first recall that, with $\mathbf{x}=(x_{1},\dots,x_{n})$ the Jack polynomials $P_{\kappa}^{(\alpha)}(\mathbf{x})$ are symmetric polynomials of $n$ variables [42, Ch. VI.6], [23, Ch. 12], [41, Ch. 7]. They are dependent on a parameter $\alpha>0$ and labelled by a partition $\kappa:=(\kappa_{1},\dots,\kappa_{n})$; in the latter each part $\kappa_{i}\in\mathbb{Z}_{\geq 0}$ is ordered as $\kappa_{1}\geq\kappa_{2}\geq\cdots\geq\kappa_{n}$. These polynomials are eigenfunctions of the differential operator
@@
$P_{\kappa}^{(\alpha)}(\mathbf{x})\mapsto\alpha^{|\kappa|}P_{\kappa}^{(\alpha)}(\mathbf{x})/h_{\kappa}^{\prime}.$
Here the precise definition of $h_{\kappa}^{\prime}$ (which depends too on $\alpha$) can be found in e.g. [23, Eq. (12.46)]. Playing the role of the Pochhammer symbol is the quantity
@@
see e.g. [23, Eq. (12.46)]. With this notation we then have as the series definition of the generalised hypergeometric function in question (see e.g. [23, §13.1])
@@
(see too [23, Eq. (13.27)] and [27, Eq. (4.32)]), where $c\in\mathbb{Z}_{\geq 1}$ (this restriction can be removed by suitably deforming the implied circular contours in (2) to Hankel loops [24, Proof of Prop. 2]) and
@@
for a derivation of this result, and its generalisation for $\alpha=1/2$ and $\alpha=2$, based on the theory of zonal polynomials, see [30]. Extending this, for any $q\in\mathbb{Z}_{\geq 0}$ we read off that
@@
For the $\beta$-generalisation of the eigenvalue PDF for complex Wishart matrices one has the functional form proportional to
$\prod_{l=1}^{N}x_{l}^{a}e^{-\beta x_{l}/2}\mathbbm{1}_{x_{l}>o}\prod_{1\leq j<k\leq N}|x_{k}-x_{j}|^{\beta};$ (2.12)
see e.g. [23, Eq. (3.16) with $a$ appropriately identified]. For this ensemble, let $E_{N,\beta}((0,s);a)$ denote the probability of there being no eigenvalues in $(0,s)$. Generalising the equality implied by (2.1) and (2.8), it was shown in [22] that for $a\in\mathbb{Z}_{\geq 0}$
$E_{\beta}^{\rm hard}((0,s);a):=\lim_{N\to\infty}E_{N,\beta}((0,s/(4N));a)=e^{-\beta s}{\vphantom{F}}_{0}^{\mathstrut}F_{1}^{(\beta/2)}(2a/\beta;(s/4)^{a}).$ (2.13)
This then relates to (2.11) with $\beta\mapsto 4/\beta$, $l=a$, $q=2/\beta-1$.
@@ 3 Matrix differential equation characterisation
The significance of this is that for $\mathbf{x}=(t)^{n}$, since the 2022 work [28] we have available a matrix differential equation characterisation of ${\vphantom{F}}_{1}^{\mathstrut}F_{1}^{(\alpha)}(a;b;\mathbf{x})$. This comes about in the setting of the integral representation [23, Cor. 2.1]
@@
[…]
$H_{p}(x)={1\over C_{p}^{n}}\int_{0}^{1}dt_{1}\cdots\int_{0}^{1}dt_{n}\,\prod_{l=1}^{n}t_{l}^{a}(1-t_{l})^{b-1}e^{-x\sum_{l=1}^{n}t_{l}}\\ \times\prod_{1\leq j<k\leq n}|t_{k}-t_{j}|^{\beta}e_{p}(1-t_{1},\dots,1-t_{n}),$ (3.3)
@@
[28, Cor. 2.5] The multidimesional integrals $\{H_{p}(x)\}_{p=0}^{n}$ satisfy the differential-difference system
$(n-p)xH_{p+1}(x)=\Big((n-p)x+B_{p}\Big)H_{p}(x)+x{d\over dx}H_{p}(x)-D_{p}H_{p-1}(x),\quad(p=0,\dots,n),$ (3.4)
@@
$B_{p}=p\Big(a+b+1+(\beta/2)(2n-p-1)\Big),\qquad D_{p}=p\Big((\beta/2)(n-p)+b\Big).$
@@
Introduce the column vector $\mathbf{H}(x)=[H_{p}(x)]_{p=0}^{n}$, and the bi-diagonal matrices
@@
$\mathbf{Y}$ $=-{\rm diag}\,[B_{0},B_{1},\dots,B_{n}]+{\rm diag}^{-}\,[D_{1},D_{2},\dots,D_{n}].$ (3.5)
@@
$x{d\over dx}\mathbf{H}(x)=(x\mathbf{X}+\mathbf{Y})\mathbf{H}(x).$ (3.6)
As a consequence [28, Remark 4.19], introducing the vector of power series solutions
$\mathbf{H}(x)=\Big[\sum_{k=0}^{\infty}c_{p,k}x^{k}\Big]_{p=0}^{n},$
@@
$c_{0,0}=S_{n}(a,b-1,\beta),\quad c_{n,0}=S_{n}(a,b,\beta),\quad c_{n-p,0}=c_{n,0}\prod_{s=1}^{p}{B_{n+1-s}\over D_{n+1-s}},\>\>(p=1,\dots,n)$
@@
We now want to make use of (3.1) in Proposition 3.1, in view of both the first and last components of $\mathbf{H}(x)$ relating to $_{1}F^{(2/\beta)}_{1}$.
@@
Consider the differential-difference system [^1] Here and in the remainder of the paper the functions $\{I_{p}(x)\}$ are not to be confused with the I-Bessel function as appearing in (1.4) and (1.7)
$(n-p)xI_{p+1}(x)=B_{p}I_{p}(x)+x{d\over dx}I_{p}(x)-pI_{p-1}(x),\quad(p=0,\dots,n),$ (3.8)
where $B_{p}$ is as in (3.4). Introduce the column vector ${\mathbf{I}}(x)=[I_{p}(x)]_{p=0}^{n}$, and the matrices
${\mathbf{W}}={\rm diag}^{+}\,[n,n-1,\dots,1],\quad{\mathbf{Z}}=-{\rm diag}\,[B_{0},B_{1},\dots,B_{n}]+{\rm diag}^{-}\,[1,2,\dots,n].$ (3.9)
@@
$x{d\over dx}\mathbf{I}(x)=(x\mathbf{W}+\mathbf{Z})\mathbf{I}(x).$ (3.10)
@@
$I_{0}(x)$ $={\vphantom{F}}_{0}^{\mathstrut}F_{1}^{(2/\beta)}\Big(\beta(n-1)+a+b+1;(x)^{n}\Big),$ (3.12)
$I_{n}(x)$ $=d_{n,0}\,{\vphantom{F}}_{0}^{\mathstrut}F_{1}^{(2/\beta)}\Big(\beta(n-1)+a+b+2;(x)^{n}\Big),$ (3.13)
and, provided all the $B_{s}$ are nonzero,
$d_{p,0}={p!\over\prod_{s=1}^{p}B_{s}},\>\>(p=1,\dots,n).$ (3.14)
@@
$(k{\mathbf{I}}_{n+1}-\mathbf{Z})\mathbf{d}_{k}=\mathbf{W}\mathbf{d}_{k-1},\>\>(k=1,2,\dots).$ (3.15)
@@
To deduce (3.8) and (3.10), we substitute in (3.4) and (3.6), respectively, $x\mapsto-x/a$, $H_{p}(-x/a)=b^{p}I_{p}(x)$ and take the limit $a\to\infty$ with $a+b$ fixed. The evaluations (3.13) follow from the evaluations of $H_{n}(x)$, $H_{0}(x)$ in terms of $\vphantom{F}{}_{1}^{\mathstrut}F_{1}^{(2/\beta)}$ noted in the sentence above Proposition 3.1. The evaluation (3.14) can be calculated as the eigenvector of $\mathbf{Z}$} corresponding to the eigenvalue 0, normalised to have its first component equal to unity. The vector recurrence relation (3.15) follows by substituting (3.11) in (3.10) and equating like powers of $x$. ∎
Analogous to the situation with (3.7), since ${\mathbf{d}}_{0}$ is known from (3.14) the vector recurrence (3.17) uniquely determines ${\mathbf{d}}_{1},{\mathbf{d}}_{2},\dots$ provided $(l{\mathbf{I}}_{n+1}-\mathbf{Z})$ is invertible for all $l\in\mathbb{Z}^{+}$. Since $\mathbf{Z}$ is lower triangular, the criterium for the latter is simply $B_{p}\notin\mathbb{Z}^{+}$, for each $p=1,\dots,n$. Relevant to (2.11) are the choices in Proposition 3.2 of $n=l$, and $a+b$ such that
$B_{p}=p\Big(q+1+(\beta/2)(l-p)\Big),$ (3.16)
@@
$\mathbf{d}_{k}=(k{\mathbf{I}}_{l+1}-\mathbf{Z})^{-1}\mathbf{W}\mathbf{d}_{k-1},\>\>(k=1,2,\dots)$ (3.17)
@@
$\mathbf{R}_{N,l}:=(N{\mathbf{I}}_{l+1}-\mathbf{Z})^{-1}{\mathbf{W}}((N-1){\mathbf{I}}_{l+1}-\mathbf{Z})^{-1}{\mathbf{W}}\cdots({\mathbf{I}}_{l+1}-\mathbf{Z})^{-1}{\mathbf{W}}.$
@@
1. Suppose we replace $q$ by q1q-1 in (3.16). Then setting $q=0$ we observe that the column vector $\delta_{p,l}]_{p=0}^{l}$ is an eigenvector of $\mathbf{Z}$} with $n=l$, allowing for the choice of initial vector corresponding to $d_{p,0}=\delta_{p,l}$ ($p=0,\dots,l$). Using this initial vector, an alternative formula to (3.18) is
@@
where the notation $(\cdot)_{l+1}$ denotes the $(l+1)$-th entry (final entry) in the corresponding column vector. In the following section, when referring to $I_{n}(x)$ in the case $q=0$ as it relates to (2.8), it will be assumed that the initial vector has been similarly modified (and thus in (3.13) $d_{n,0}$ is replaced by unity — without this modification (3.13) is invalid since $d_{n,0}$ as specified by (3.14) involves a division by zero).
2. For efficient computation using the matrix differential equation, it is advantageous to forgo (3.18) in favour of a recursive computation based on (3.15). See the second subsection of Appendix B for a complexity analysis when using this method to compute $\{T_{l}(N)\}$ in (1.3), which moreover is compared against the formalism used in [29, 12] for the same purpose.